(4 pts.) Given the following demand functions find the elasticity of demand at a price of $9 (give your answer as a fraction and show your work on a separate page): 2. a. Qp 180- 2P2 ED Circle one: Unit elasitic Elastic Inelastic b. Qp 45 6PO.5 ED Circle one: Unit elasitic Elastic Inelastic
Answer #1

Price elasticity = dQ/dP x P/Q

a) PE = -4P x P/Q = - 4 x 92/(180 - 2 x 92) = - 18 (Elastic demand, absolute value > 1)

b) PE = -3/P0.5 x P/(45 - 6P0.5) = -3/3 x 9/(45 - 6 x 3) = - 0.33 (Inelastic demand, absolute value < 1)